If you buy 12 lb of Possum meat for $2/lb, then 20 lb for $1.50/lb & 7 lb for $4/lb,the average price you paid was $2.43/lb, right?

Answer by JulietG(1812)

12# @ $2/lb = $24 and 12#
20# @ 1.50/lb = $30 and 20#
7# @ $4/lb = $28 and 7#
Totals
39# for $82
Divide 82 by 39 to get $2.10/lb